### v3

# The DeepSpeak Dataset

Abstract
We describe a large-scale dataset – DeepSpeak – of real and deepfake footage of people talking and gesturing in front of their webcams. The real videos in this dataset consist of a total of 50 hours of footage from 500 diverse individuals. Constituting more than 50 hours of footage, the fake videos consist of a range of different state-of-the-art avatar, face-swap, and lip-sync deepfakes with natural and AI-generated voices. We are regularly releasing updated versions of this dataset with the latest deepfake technologies. This preprint describes the construction of versions 1.0, 1.1, and 2.0. This dataset is made freely available for research and non-commercial uses; requests for commercial use will be considered.
Keywords Deepfakes  $\cdot$ Generative AI  $\cdot$ Dataset  $\cdot$ Digital Forensics
1 Introduction
Large-scale and diverse datasets are critical to the academic literature to aid in both system development and evaluation. We describe an audio and video dataset designed to aid the digital-forensic community [^1] We named our dataset DeepSpeak before the release of the DeepSeek AI model. This dataset consists of 100 hours of real and deepfake video of people talking and gesturing. The real videos were recorded with consent from the participants. The deepfakes consist of avatar (from three generators), face-swap deepfakes (with multiple variants from three generators), lip-sync deepfakes (from four generators), and audio deepfakes (from three generators) spliced into a subset of the lip-sync deepfakes. Shown in Table 1 is a comparison of the characteristics of the different versions of our dataset as compared to recent (past seven years) datasets.
We describe the details of the collection of the real video and how the various deepfake audio and video are created. Each version of the dataset is available by request at the following locations:
Given the pace at which the underlying deepfake technology is progressing, it is critical that evaluation datasets keep up with the latest technologies. To this end, we expect to release an updated dataset once to twice a year, expanding all aspects of the dataset.
Release Unique Original Deepfake
Name Year Identities Footage Consent Faceswap Lipsync Avatar Audio Footage
FaceForensics  [25] 2018 NA 1,004 N $\checkmark$ - - - 2,008
FaceForensics++ [26] 2019 NA 1,000 N $\checkmark$ - $\checkmark$ - 4,000
DFDC [6] 2020 3,426 23,654 Y $\checkmark$ - - - 104,500
Celeb-DF [15] 2020 59 590 N $\checkmark$ - - - 5,639
WildDeepfake [43] 2020 707 3,805 N NA NA NA NA 3,509
DeeperForensics [11] 2020 100 50,000 Y $\checkmark$ - - - 10,000
FakeAVCeleb [12] 2021 600 570 N $\checkmark$ $\checkmark$ - - 25,000
ForgeryNet [10] 2021 5,400 99,630 Y/N $\checkmark$ - - - 121,617
LAV-DF [2] 2022 153 36,431 N - - $\checkmark$ - 99,873
DF40 [35] 2024 NA NA N - 100,000+
DeepSpeak v1.0 2024 219 6,226 Y $\checkmark$ $\checkmark$ - $\checkmark$ 6,799
DeepSpeak v1.1 2025 220 6,667 Y $\checkmark$ $\checkmark$ - $\checkmark$ 6,796
DeepSpeak v2.0 2025 280 9,376 Y $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 7,209
Table 1: A comparison of forensic-themed public datasets. Although not the most informative metric, we report original and deepfake footage as number of videos for consistency with previous published datasets (NA: not available).
2 Data Collection
Participants were crowd-sourced through the Prolific research recruitment platform. Participants were asked to give their consent for including their recordings, without any other identifying information, in a public dataset [^2] The precise consent language was: “This dataset will be used for research purposes for detecting deepfakes. Please note that your recordings will be made public in a dataset, but no other identifying information will be shared outside of our research group.”. Participants were selected from a stratified sample ensuring equal distribution of gender, with all participants being native English speakers and U.S. residents.
Each participant completed a survey in which they were instructed to record themselves responding to 32 (version 1.0) and 35 (version 2.0) separate prompts (see Appendix A). Participants were paid \$7 for their time. The first two prompts were used for voice-clone training data (see Section 3). The remaining prompts were divided into four categories: (1) standardized scripted responses in which each participant read the same prompt; (2) randomized scripted responses in which participants read a randomized prompt; (3) unscripted responses in which participants responded to questions; and (4) actions in which participants performed simple actions.
Both audio and video were recorded using a custom-built Google Chrome web application. Recordings were captured as .webm files with a bitrate of 8 megabits per second, using the Google VP9 codec for video compression. The target resolution was set at 1280 $\times$× 720 pixels, but users with limited bandwidth were able to record at lower resolution of 640 $\times$× 480 pixels. The JavaScript and Python repository for this web application is available at https://github.com/sbarrington/UCB-web-video-recorder-app-main.
Participants were given written and visual instructions to allow them to practice recording themselves and test their hardware. Participants were asked to adhere to a series of recording conditions intended to improve consistency within the overall dataset, including positioning themselves centrally within their web camera frame, sitting in a well lit room, not allowing other faces or people to be present in the frame, and minimizing background noise.
For scripted responses (see Appendix A), participants were asked to record themselves repeating a short script while looking into the camera. Scripted responses were obtained using transcripts of the TIMIT dataset [8]. The TIMIT dataset consists of 462 real female and male American-English speakers, uttering a total of 1,718 short-to-medium length phonetically-rich sentences. Sentences of length less than the mean of 50 characters were removed. Ten sentences were then selected at random for the standardized scripts read by all participants. The remaining 728 sentences comprised the randomized scripts, for which each participant read a random sample of 10.
Participants then responded to four open-ended unscripted questions and asked to aim for a response that was close to 30 seconds in length. These were followed by quick-fire questions in which they repeated the question and provided a short response. Following the scripted and unscripted responses, participants were asked to perform simple actions using head and hand gestures. See Appendix A for the full list of prompts and scripts used.
We manually removed any invalid responses from the final dataset. This included ensuring that only a single person was visible in the video, the scene was reasonably well lit, and each submitted video clip contained a valid audio and visual track.
2.1 Version 1.0
We collected data from 220 participants for version 1.1 (version 1.0 had only 219 participants due to a corrupt file). The demographics of the participants are as follows (some participants identified with more than one race/ethnicity):
  • Age: Range = 18-75 years, Mean = 38 years; standard deviation = 11.5 years
  • Gender: 110 male, 106 female, 4 non-binary
  • Race/Ethnicity: 159 White/Caucasian, 38 Black/African American, 25 Asian, 3 American Indian/Alaska Native, 1 Native Hawaiian/Other Pacific Islander, 5 other.
Participants were asked to perform four simple visual actions: (1) wave their hand in front of their face while counting 1,2,3; (2) look down and right, straight down, and down and left, each time holding and counting 1,2,3; (3) look up and right, straight up, and up and left, each time holding and counting 1,2,3; (4) lean towards the camera without counting.
The audio/video recordings were converted from their initial .webm format to .mp4 and organized by prompt type using the H.264 codec with the libx264 library (this final video conversion is not performed in version 2.0 in order to minimize re-encoding).
Because version 1.1 corrects errors from version 1.0, we recommend using this version.
read wave hand turn head right/up tilt head down lean in
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: Representative examples of different prompt types (see Appendix A).
2.2 Version 2.0
We collected data from 280 participants for version 2.0. The demographics of the participants were as follows (some participants identified with more than one race/ethnicity):
  • Age: Range = 18-73 years, Mean = 38 years; standard deviation = 11.6 years
  • Gender: 146 male, 129 female, 3 non-binary, 2 not provided
  • Race/Ethnicity: 203 White/Caucasian, 49 Black/African American, 20 Asian, 11 American Indian/Alaska Native, 1 Native Hawaiian/Other Pacific Islander, 10 other, 1 prefer not to say.
Participants were asked to perform the same four simple visual actions as in version 1.0: (1) wave their hand in front of their face while counting 1,2,3; (2) look down and right, straight down, and down and left, each time holding and counting 1,2,3; (3) look up and right, straight up, and up and left, each time holding and counting 1,2,3; (4) lean towards the camera without counting). An additional set of three new actions were added to this version: (5) pretending to yawn; (6) pretending to laugh for between 2 to 5 seconds; (7) clapping loudly three times, pausing for about 5 seconds between claps.
While audio/video recordings for version 1.0 were converted from their initial .webm format to .mp4 using the H.264 codec with the libx264 library, an FFmpeg “copy” command was used in version 2.0 without re-encoding to preserve recording quality.
3 Audio Generation
Participants were first asked to record themselves reading 10 consecutive phonetically-rich sentences, sourced from List 1 of the standard Harvard Sentences [27]. Participants were then asked to repeat the standard elicitation paragraph from the Speech Accent Archive, a phonetically comprehensive passage comprising a breadth of vowels and consonants [33]. These two scripted responses were used for the purpose of voice cloning, and had an average length of 30 seconds.
Using each participant’s cloned voice, a synthetic audio was created in their voice saying the same thing as in the original audio/video. For the unscripted responses, the original audio was transcribed using OpenAI’s Whisper, and for the scripted responses, we assumed that the participant correctly read the script. These text transcriptions were then provided to each voice cloning generators’ API to generate matching synthetic voices. The following generators were used for each version of the DeepSpeak dataset.
3.1 Version 1.0
Voice clones were generated using the Elevenlabs Create Voice endpoint [^3] https://elevenlabs.io/docs/api-reference/voices/add. The Create Speech API [^4] https://elevenlabs.io/docs/api-reference/text-to-speech/convert was used for speech synthesis. The eleven_multilingual_v2 model was used, with audio output returned in MP3 format at 44.1 kHz. The API was accessed in April 2024.
3.2 Version 2.0
As in version 1.0, voice clones in version 2.0 used the same two scripted prompts for all voice-cloning generators.
The ElevenLabs API was accessed as described in the previous section.
Voice clones were generated using the PlayAI Instant Voice Cloning endpoint [^5] https://docs.play.ht/reference/api-create-instant-voice-clone. The pyht Python package, with TTSOptions set to default parameters was used for speech synthesis, with audio output provided in MP3 format, at a sampling rate of 24 kHz.
Voice clones were generated using the Speechify v1 Voices API [^6] https://docs.sws.speechify.com/v1/api-reference/api-reference/tts/voices/create. The Speech API endpoint [^7] https://docs.sws.speechify.com/v1/api-reference/api-reference/tts/audio/speech was used for speech synthesis. The simba-english model was used, with output audio returned in WAV format at 48 kHz.
All API calls were made in November 2024.
4 Video Generation
We generated three types of video deepfakes: face swap, lip sync, and avatar, each of which is described next. Between version 1.0 and 2.0 of DeepSpeak, deepfake generators improved considerably. Because we are primarily interested in creating a dataset using state-of-the-art generators, we do not, in version 2.0, re-use the same generators as in version 1.0.
All source videos were re-encoded using FFmpeg with the H.264 codec at a constant rate factor of 18. The encoding present was set to slow, and the original audio stream was preserved without re-encoding.
4.1 Face-Swap
Face-swap deepfakes are created by replacing – eyebrow to chin and cheek to cheek – the original identity in a video with a new identity. In order to create a visually compelling face-swap deepfake, each identity in the full dataset is paired with a perceptually similar face as follows. Each identity is first represented by a mean CLIP embedding [22] of five randomly extracted video frames. For each identity a unique matched identity is assigned using the agglomerative clustering algorithm with cosine distance and cluster size constraint from the scikit-learn library [^8] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.AgglomerativeClustering.html. Representative examples of these pairs are shown in Figures 2 and 3.
After pairing each identity, the video of the original identity and a single frame of the matched identity are provided to the face-swap synthesis engine. The single frame is initially chosen to be the fifth frame in a randomly selected video of the matched identity. We found that if the eyes are closed in the matched face, the resulting face-swap deepfake suffered in quality. As such, we used MediaPipe [18] to extract facial features and ensured that the distance between the top and bottom eyelid landmarks was greater than a specified threshold. If this constraint failed, then the tenth frame was selected for consideration; this process was repeated, skipping five frames each time, until a suitable frame was found.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Four representative examples of matched identities from version 1.0.
4.1.1 Version 1.0
We used three variants of the open-source library FaceFusion as follows:
  • FaceFusion The first face-swap configuration invokes the FaceFusion [28] library with its default parameters. FaceFusion, built on InsightFace [24], localizes the face in the original video, performs 3D landmark estimation of the face, and superimposes the appropriately transformed matched face. This is followed by a set of post-processing steps to hide edge artifacts and improve temporal consistency.
  • FaceFusion + GAN: The second face-swap configuration extends the above FaceFusion configuration by enhancing each generated frame using the CodeFormer GAN [41]. This model corrects rendering discrepancies–mostly misshapen teeth and lips–and increases the overall photorealism of the face.
  • FaceFusion Live: The third face-swap configuration wraps the same FaceFusion configuration in a simulated live input streaming environment. Because our hardware (a Tesla T4 GPU) cannot generate video frames in real-time ($2424$ to $3030$ frames per second (fps)), the effective frame-rate of the generated video is decreased to approximately $1212$ fps.
4.1.2 Version 2.0
Between the construction of version 1.0 and version 2.0, FaceFusion expanded to include a larger suite of face-swap generators. For version 2.0, we used four variants from the FaceFusion library:
  • INSwapper: The first face-swap configuration invokes the INSwapper-128 model [30] with default parameters. This model leverages InsightFace [24] to localize 3D facial landmarks and superimpose a transformed face of the matched identity onto the source identity video. This version is most similar to FaceFusion in version 1.0.
  • INSwapper + CodeFormer: The second face-swap configuration extends the above INSwapper configuration by enhancing each frame of the generated deepfake using CodeFormer [40]. This model was trained to fix any misshapen teeth or lips and improve the overall photorealism of the generated face. This version is most similar to FaceFusion+GAN in version 1.0.
  • SimSwap: The third face-swap configuration invokes the SimSwap-256 model [3] with default parameters. Unlike INSwapper, which directly modifies the target face based on detected facial landmarks, SimSwap extracts a latent identity representation to guide the generation of new frames.
  • SimSwap + RestoreFormer: The fourth face-swap configuration extends the above SimSwap configuration by enhancing each frame of the generated deepfake using RestoreFormer++ [32]. This model, trained on degraded photographs, corrects any structural deficiencies and enhances the photorealism of the generated frames.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Four representative examples of matched identities from version 2.0.
4.2 Lip-Sync
Whereas the face-swap deepfake (Section 4.1) replaces an entire face with a new identity, a lip-sync deepfake modifies the mouth region to be consistent with a different audio track.
Given an original video and associated audio, we create two types of lip-sync deepfakes: (1) a lip-sync deepfake with an audio of the same identity extracted from a different video (i.e., the audio and video are now mismatched); and (2) a lip-sync deepfake with an AI-generated voice of the same identity (Section 3) with a transcript taken from a different video.
4.2.1 Version 1.0
Two types of lip-sync deepfakes were created using the following models:
  • Wav2Lip: The first lip-sync configuration invokes the Wav2Lip model [21] with its default parameters. Wav2Lip is a neural network trained in a GAN-like generator-discriminator fashion. The generator architecture is a LipGAN [13] trained on the LRS2 dataset [1].
  • VideoRetalking: The second lip-sync configuration invokes the VideoRetalking pipeline [4] with its default parameters. The input video is passed through three neural networks: (1) a semantic-guided reenactment model, which stabilizes the expression in the video; (2) a lip-sync model, which renders a new mouth and chin area matching the audio; and (3) a face enhancer, which fixes rendering discrepancies. The specific models employed in the first two stages are L-Net and D-Net; the third model is the CodeFormer GAN [41].
4.2.2 Version 2.0
Two types of lip-sync deepfakes were created using the following models:
  • Diff2Lip: The first lip-sync configuration invokes the Diff2Lip model [19] with default parameters. After localizing the face in the video, Diff2Lip adds noise over the mouth region and proceeds as an audio-conditioned diffusion model. It was trained on the Voxceleb2 [5] and LRW [36] datasets.
  • LatentSync: The second lip-sync configuration invokes the LatentSync model [14] with default parameters. While LatentSync is, at its core, a diffusion model similar to Diff2Lip, it differs in two aspects. First, it encodes the audio constraint as Whisper embeddings [23] instead of chunks of audio signal. Second, it uses a noise mask delineated by facial landmarks to exactly match the shape of the face as opposed to a rectangular bounding box. LatentSync was trained on the VoxCeleb2 [5] and HDTF [38] datasets.
4.3 Avatar
Avatar deepfakes animate the head and lip movements of a static image to match a target video or audio track. Unlike face-swap and lip-sync deepfakes, which modify an existing video, avatar deepfakes generate movement from a single static image. Version 1.0 contains no avatar deepfakes because these models were still in their infancy.
4.3.1 Version 2.0
Three types of avatar deepfakes were created using the following models:
  • LivePortrait: The first avatar configuration invokes the LivePortrait model [9] with its default parameters. LivePortrait is a multi-stage model that first extracts the identity and motion embeddings, which are warped into a joint embedding that is later decoded into pixel space. This model was trained on the VoxCeleb [20], MEAD [31], RAVDESS [17], and AAHQ [16] datasets.
  • HelloMeme: The second avatar configuration invokes the HelloMeme [37] model with its default parameters. HelloMeme consists of three modules: a reference module, which extracts an identity embedding; a control module, which extracts information about the head and mouth shape; and a diffusion module, which generates the resulting video frames. This model was trained on the CelebV-HQ [34] and VFHQ [34] datasets.
  • Memo: The third avatar configuration invokes the Memo model [39] with its default parameters. Memo is a diffusion model that combines identity, voice, and emotion embeddings as diffusion constraints. It was trained on a compilation of the HDTF [38], VFHQ [34], CelebV-HQ [42], MultiTalk [29], and MEAD [31] datasets.
LivePortrait and HelloMeme take as input a single image of a person to animate with a video (and associated audio) that drives this animation. For these two generators, the avatar deepfakes contain only real audio from the original driving video. Memo takes as input a single image of a person to animate with only an audio that drives this animation. In this case, the audio can be either real or fake.
4.4 Summary
Shown in Table 2 is a breakdown of the total size in file size (GB), file counts (N), and video length (hrs). Also shown is the breakdown for a 80/20 training/testing split. Across these splits, there is no overlap in facial or voice identities.
Total Train Test
ver size (GB) size (N) size (hrs) real (N [hrs]) fake (N [hrs]) real (N [hrs]) fake (N [hrs])
1.0 40 13,025 44.3 4,902 [13.9] 5,300 [21.0] 1,324 [3.7] 1,499 [5.8]
1.1 46 13,463 48.0 5,251 [16.8] 5,299 [21.0] 1,416 [4.4] 1,497 [5.8]
2.0 124 16,585 52.7 7,513 [23.6] 5,793 [18.6] 1,863 [5.8] 1,416 [4.6]
Table 2: A breakdown of the total size (gigabytes (GB), number of files (N), and length in hours (hrs)) of each version of the DeepSpeak dataset.
5 Discussion
Our primary motivation for creating this dataset is to support the media forensics research community and the development and refinement of techniques to detect deepfake audio, image, and video. At the same time, we can envision other uses cases and will consider other academic and commercial requests for access to the DeepSpeak dataset.
The world of generative AI and media forensics is fast moving. It is, therefore, important that shared datasets be regularly updated to keep up with the latest trends. To this end, we expect to release updates to this dataset once to twice a year. To help us serve the community better, we welcome feedback, comments, requests for future releases of this dataset. An online feedback form is available at https://huggingface.co/datasets/faridlab/deepspeak_v1.
Acknowledgments
This work was partially supported by Google/YouTube and the University of California Noyce Initiative. We are grateful to ElevenLabs (https://elevenlabs.io) and PlayAI (https://play.ht/) for granting us API access for voice generation.
Appendix A Prompts
Video Action Prompts
A.1 Video Action Prompts
Versions 1.0 and 2.0
Lean forwards towards the camera and back to your starting position.
1. Look down and to the right, and slowly count out loud to three.
2. Look down and to the middle, and slowly count out loud to three.
3. Look down and to the left, and slowly count out loud to three.
1. Look up and to the left, and slowly count out loud to three.
2. Look up and to the middle, and slowly count out loud to three.
3. Look up and to the right, and slowly count out loud to three.
Wave your hand back and forth across your face four times while counting out loud.
Read each question aloud, followed immediately by your answer.
1. What is my favorite food?
2. What is my favorite movie?
3. What did I have for breakfast?
4. Where would I most like to travel and why?
5. If I could time travel, where in the past or future would I like to go?
Version 2.0 only
Pretend to yawn, perhaps covering your mouth with your hand.
Pretend to laugh for between 2 to 5 seconds.
Clap loudly three times, pausing for about 5 seconds between claps. Please make sure your clap makes a loud sound.
Table 3: Video action prompts.
A.2 Standardized Scripted Prompts
We apply auditory modeling to computer speech recognition.
He sank back sighing and was soon asleep again.
The mango and the papaya are in a bowl.
Will you tell me why.
That experience holds a lesson for us all in regard to birth control today.
That is what childhood is he told himself.
As a rule part time farmers hire little help.
The new birth is miraculous and mysterious.
The fear of punishment just didn’t bother him.
The figure in the corner belched loudly a deep liquid eruption.
Table 4: Standardized scripted prompts, consisting of ten separate sentences for version 1.0 and 2.0.
A.3 Voice Cloning Prompts
The birch canoe slid on the smooth planks.
Glue the sheet to the dark blue background.
It’s easy to tell the depth of a well.
These days a chicken leg is a rare dish.
Rice is often served in round bowls.
The juice of lemons makes fine punch.
The box was thrown beside the parked truck.
The hogs were fed chopped corn and garbage.
Four hours of steady work faced us.
A large size in stockings is hard to sell.
Please call Stella. Ask her to bring these things with her from the store: Six spoons of fresh snow peas, five thick slabs of blue cheese, and maybe a snack for her brother Bob. We also need a small plastic snake and a big toy frog for the kids. She can scoop these things into three red bags, and we will go meet her Wednesday at the train station.
Table 5: Voice cloning input prompts for version 1.0 and 2.0, consisting of ten consecutive sentences and one continuous paragraph.
A.4 Randomized Scripted Prompts
Table 6: 728 randomized scripted prompts for version 1.0 and 2.0, from which each participant only responded to 10.
She had your dark suit in greasy wash water all year. Her classical performance gained critical acclaim. The diagnosis was discouraging however he was not overly worried.
Rich looked for spotted hyenas and jaguars on the safari. The football team coach has a watch thin as a dime. He found an empty bench opened a newspaper and stretched his legs before him.
The sermon emphasized the need for affirmative action. You’re so preoccupied that you’ve let your faith grow dim. Gus saw pine trees and redwoods on his walk through sequoia national forest.
Approach your interview with statuesque composure. The annoying raccoons slipped into phil’s garden every night. Some tore entirely through the whipsawed post oak.
Markets should become more competitive as consumers become more selective. Pledge to participate in nevada’s aquatic competition. The advertising verse of plymouth variety store never changes.
The eastern coast is a place for pure pleasure and excitement. She always could sense the shag end of a woolly day. These planets were much bigger nearly all capable of holding an atmosphere.
He had not covered a hundred yards before a gun crashed from somewhere behind. Between meetings he helps the president keep track of delegated matters. A crab challenged me but a quick stab vanquished him.
Just why anybody should wish to start a riot the executive officer didn’t know. They used an aggressive policeman to flag thoughtless motorists. The source is known so there is no necessity to remove insecticide residues.
A toothpaste tube should be squeezed from the bottom. A precision transit is set up so that it is lined with respect to true north. Two gas lamps were no more than a misleading glow.
Those who are not purists used canned vegetables when making stew. Unless we send out the whole pie their pieces mean nothing. The easygoing zoologist relaxed throughout the voyage.
His failure to open the store by eight cost him his job. Smash lightbulbs and their cash value will diminish to nothing. Etiquette mandates compliance with existing regulations.
Remember to allow identical twins to enter freely. She always jokes about too much garlic in his food. The government sought authorization of his citizenship.
Children can consume many fruit candies in one sitting. Bob bandaged both wounds with the skill of a doctor. It gives social guidance and direction and makes for programs of social action.
No signs of these no gross hemorrhage of lungs heart brain or stomach. Herb’s birthday occurs frequently on thanksgiving. All about him stood tombstones his own sensitive great hands had fashioned.
With the spring rains the flow rose rapidly due to infiltration in open sewers. Scholastic aptitude is judged by standardized tests. We could barely see the fjords through the snow flurries.
A concurrent effort is needed to make oceanographic data useful on the spot. This may be of overriding importance in considering military objectives. A moth zig zagged along the path through otto’s garden.
Buying a thoroughbred horse requires intuition and expertise. His technique is ample and his musical ideas are projected beautifully. Al received a joint appointment in the biology and the engineering departments.
The high security prison was surrounded by barbed wire. Count the number of teaspoons of soysauce that you add. While waiting for chipper she crisscrossed the square many times.
No girl would go this far to fool a man so she could kill him. Northern liberals are the chief supporters of civil rights and of integration. That stinging vapor was caused by chloride vaporization.
Whether historically a fact or not the legend has a certain symbolic value. The library has open shelves even in the unbound periodical stockroom. Steph could barely handle the psychological trauma.
In every major cloverleaf traffic sometimes gets backed up. Davy mathews it’s disgusting the way you’re always eating. As a precaution the outlaws bought gunpowder for their stronghold.
The entire length of the street could be raked with rifle fire from this barn. Severe myopia contributed to ron’s inferiority complex. May i order a strawberry sundae after i eat dinner.
And never show my face or my truck around here again. Any organism that falters or misperceives the signals or weakens is done. Valley lodge yearly celebrates the first calf born.
Then again there’s always that lovely old pastime of hooking or braiding rugs. A young mouse scampered across the field and disappeared. What possessed you to tell me a clotheshorse would be a good idea.
Employee layoffs coincided with the company’s reorganization. The antithesis of the ecumenical and the local then no longer exists. The social and psychological consequences of this continue to affect the area.
Make a paste of brown sugar and mustard and spread lightly over scored surface. In many of his poems death comes by train a strongly evocative visual image. In most cases we recognize certain words persons animals or objects.
Cooperation along with understanding alleviate dispute. Gregory and tom chose to watch cartoons in the afternoon. Whoever cooperates in finding nan’s cameo will be rewarded.
Such legislation was clarified and extended from time to time thereafter. But none of this could soothe the exacerbated nerves. He waited until they were inside the elevator and then said now what do we do.
Coffee is grown on steep jungle like slopes in temperate zones. We like bleu cheese but victor prefers swiss cheese. My desires are simple give me one informative paragraph on the subject.
Alice i tell you my feet are like chopped beefsteak. Almost everybody in the senior class is married students say dogmatically. They all agree that the essay is barely intelligible.
It moved in a silver arc toward his throat then veered downward. His legs pumped furiously his long black hair streamed out behind him. If we left one we’d have to wipe it for fingerprints.
The rich should invest in black zircons instead of stylish shoes. And the surface is driven back in its very surfaceness only by this contrast. Since a fall or blow might have caused it a cold pack was usually first aid.
Positive results start when it goes towards the hand you use to make your mark. The preschooler couldn’t verbalize her feelings about the emergency conditions. Are we as safe as we should be from such a disaster.
Before deriving this formula we explain what we mean by problems of this kind. My sincere wish is that he continues to add to this record he sets here today. Neither his appetites his exacerbations nor his despair were kin to yours.
The giant redwoods shimmered in the glistening sun. The cow wandered from the farmland and became lost. We have become amateur insurance experts and fine feathered yard birds.
Privately he created and magnified an image of himself as a hired assassin. Both loved the out of doors including mountain climbing and horseback riding. And their chroniclers are not the dramatic poets but the prose novelists.
According to my interpretation of the problem two lines must be perpendicular. In the course of its inquiry it took testimony from only seven witnesses. Seamstresses attach zippers with a thimble needle and thread.
Although they drew light ground fire they saw no signs of activity. Nonprofit organizations have frequent fund raisers. Boys and men go along the riverbank or to the alcoves in the top arcade.
The response of reaction is dominated by a concern for what is vanishing. The fat man has trouble buying life insurance or has to pay higher premiums. On all sides doors were being slammed in his face.
Which long article was opaque and needed clarification. To what extent such low density applies to micrometeorites is unknown. The patient and the surgeon are both recuperating from the lengthy operation.
The haunted house was a hit due to outstanding audio visual effects. Those answers will be straightforward if you think them through carefully first. This was chiefly because of the bluish white autofluorescence from the cells.
Biologists use radioactive isotopes to study microorganisms. At the left is a pair of dressy straw pumps in a light but crisp texture. He reached out and felt the bath towel hanging on the towel rack over the tub.
Thus far the advances made have been almost entirely along functional lines. There was a gigantic wasp next to irving’s big top hat. With her sharp tongue she’d have cut his pompousness to ribbons.
However when labor disputes arise its provisions come clearly into play. Confusion became chaos each succeeding day brought new acts of violence. His tough honesty condemns him to a solitary and difficult existence.
Later we shall see what happened when an emperor took this idea too literally. Thirty five they rode at a measured pace through the valley. Technical writers can abbreviate in bibliographies.
A clearly recognized exception is a statutory merger or consolidation. Hiding out like this won’t get him anything except more trouble or bullet. I gave them several choices and let them set the priorities.
The water contained too much chlorine and stung his eyes. An adult male baboon’s teeth are not suitable for eating shellfish. Begin the examination of a site with a good map and aerial photos if possible.
These exclusive documents must be locked up at all times. Butterscotch fudge goes well with vanilla ice cream. Index words and electronic switches may be reserved in the following ways.
Thus there is a clearer division of authority administrative and legislative. This is a significant advance but its import should not be exaggerated. Todd placed top priority on getting his bike fixed.
He was above all a friend seeker almost pathetic in his eagerness to be liked. So we note approvingly a fresh sample of unanimity. No she would not pretend modesty but neither must she be crudely bold.
Roy ignored the spurious data points in drawing the graph. Whosoever violates our rooftree the legend states can expect maximal sorrow. A sailboat may have a bone in her teeth one minute and lie becalmed the next.
The new suburbanites worked hard on refurbishing their older home. He played basketball there while working toward a law degree. He had accordingly cultivated eccentricity to the point of second nature.
The frightened child was gently subdued by his big brother. No manufacturer has taken the initiative in pointing out the cost involved. The desire and ability to read are important aspects of our cultural life.
A complete plan we have made limited application of the parallel ladder plan. The public is now armed with sophistication and numerous competing media. The barracuda recoiled from the serpent’s poisonous fangs.
He liked to nip ear lobes of unsuspecting visitors with his needle sharp teeth. Two miles northeast then five miles southwest that sort of thing. He picked up nine pairs of socks for each brother.
Brush fires are common in the dry underbrush of nevada. The dead spirits occupied a prominent place in every hope and in every fear. Behind him billowed a small pungent cloud of smoke.
Suburban housewives often suffer from the gab habit. Will you please confirm government policy regarding waste removal. Perhaps this is what gives the aborigine his odd air of dignity.
Thereupon followed a demonstration that tyranny knows no ideological confines. Research into several cultures has proven her position to be a mistaken one. She had jumped away from his shy touch like a cat confronted by a sidewinder.
Faces may be made into candles by filling with melted wax and a wick. In most discussions of this phenomenon the figures are substantially inflated. C’mon he whispered four levels about three feet down so don’t fall.
The sudden solitude had lost its momentary charm and become oppressive. Vital questions would be quickly answered according to a preprepared agenda. The mayan neoclassic scholar disappeared while surveying ancient ruins.
They also furnish proof that in modern war message sending must be monitored. But problems cling to pools as any pool owner knows. Personal predispositions tend to blunt the ear and in turn the voice as well.
Only rarely is attention given to accurate progress reports and evaluation. Both have excellent integration of their fiscal tax collection year calendars. Need for novelty may be a symptom of cultural fatigue and instability.
The mixing head moves back and forth slowly across the width of the receptacle. First they wanted to clarify a tantalizing bizarre enigma. Shell shock caused by shrapnel is sometimes cured through group therapy.
The groundhog clearly saw his shadow but stayed out only a moment. Two cars came over a crest their chrome and glass flashing. He really crucified him he nailed it for a yard loss.
The trouble is that like many symbols it doesn’t seem a very realistic one. He ripped down the cellophane carefully and laid three dogs on the tin foil. The flat bottomed boat swung slowly to the pull of the current.
Scientific progress comes from the development of new techniques. The cigarettes in the clay ashtray overflowed onto the oak table. The most recent geological survey found seismic activity.
Men believed they could control nature by obeying a moral code. He spoke briefly sensibly to the point and without oratorical flourishes. I’ll have a scoop of that exotic purple and turquoise sherbet.
But it must be remembered that the plan should not be oriented geographically. Come on let’s hurry down before they lock up for the day. The boston ballet overcame their funding shortage.
Yet the spirit which lives in community is not identical with the community. Wine glass heels are to be found in both high and semi heights. Naturally no woman can ever completely monopolize the sexual initiative.
The local drugstore was charged with illegally dispensing tranquilizers. Flying standby can be practical if you want to save money. Assume for example a situation where a farm has a packing shed and fields.
It made no difference that most evidence points to an opposite conclusion. Rob made hungarian goulash for dinner and gooseberry pie for dessert. Shall we flip a coin to see which of us goes first.
Jokes cartoons and cynics to the contrary mothers in law make good friends. Be careful that you keep adequate coverage but look for places to save money. But fenced or unfenced no pool side is the place for running or horseplay.
Another field had given him fame enough to satisfy any egotist. The cowboy’s humorous name for a cow givin’ milk was a milk pitcher. For wasp stings onion juice obtained by scraping an onion gave quick relief.
He may have a point in urging that decadent themes be given fewer prizes. Now don’t tell me what a good ball player you are. They know little about their machinery beyond mechanical details.
The legislature met to judge the state of public education. He is forced to play for little money and must often take another job to live. Some have walked through pain and sorrow to bring you their message of hope.
The thick elm forest was nearly overwhelmed by dutch elm disease. Needless to say my art suffered drastically during this turbulent period. Once you finish greasing your chain be sure to wash thoroughly.
When peeling an orange it is hard not to spray juice. They stayed at hotels and boardinghouses or at private homes. Internal national responsibility now a truism need not be documented.
Ironically enough in this instance such personal virtues were a luxury. Far more frequently overeating is a result of a psychological compulsion. Solid concrete blocks relatively heavy and dense are used for this shelter.
The overweight charmer could slip poison into anyone’s tea. Along the main thoroughfares hardly a house had not been peppered. We experience distress and frustration obtaining our degrees.
The lack of heat compounded the tenant’s grievances. Selecting bunks by economic comparison is usually an individual problem. Yeah seems so don’t it the boy laughed hugging her close.
He showed puny men attacked by splendidly tyrannical machines. He can for example present significant university wide issues to the senate. The moisture in my eyes is from eyedrops not from tears.
Your voice is delightful he approved with a warm smile. Differences were related to social economic and educational backgrounds. He crossed the next meadow and climbed a tree where the jungle trail resumed.
This truth that the moral law is natural has other important corollaries. If they are not ellipsoids the conclusions will be a reasonable approximation. But he was very much like his associates in his hatred of camp routine.
His blue eyes sought the shimmering sea of haze ahead. The single curve line represents a specific formulation in a test example. He saw a pint sized man with a graying spade beard and an unusually large head.
The word means it won’t boil away easily nothing else. These air or gas bubbles make highly functional thermal barriers. The proof that you are seeking is not available in books.
A screwdriver is made from vodka and orange juice. Unit prices for state vehicles are invariably lower than to the general public. Every single problem touched on thus far is related to good marketing planning.
While one element is announcing progress another is delineating its problems. A covered container such as a kitchen garbage pail might do as a toilet. Why the hell didn’t you come out when you saw them gang up on me.
Many wealthy tycoons splurged and bought both a yacht and a schooner. Chip postponed alimony payments until the latest possible date. A concept of responsibility is in process of articulation and establishment.
A domestic automatic washer that will give equivalent results may be used. The compounds are divided according to composition into seven categories. Draw every outer line first then fill in the interior.
Usually they titter loudly after they have passed by. Chocolate and roses never fail as a romantic gift. It was as blissful and fulfilling a night as any bride ever experienced.
Yes indeed we too can see a warlike host of infidels encamped against us. Lifting her skirts she climbed in never relinquishing her grip on his arm. Again these blocks were set in resin saturated glass cloth and nailed.
Often they are able to get in only because the area is declining economically. His prescription hot and cold compresses to increase her absorption of water. Some make beautiful chairs cabinets chests doll houses etc.
For girls the overprotection is far more pervasive. When mold has more than one design cavity make individual paper patterns. One upmanship is practiced by both sides in a total war.
The shot reverberated in diminishing whiplashes of sound. Some observers speculated that this might be his revenge on his home town. The tragic stage is a platform extending precariously between heaven and hell.
Bruises and black eyes were relieved by application of raw beefsteak. She knew she was feeling afraid and inwardly laughed at herself. However this inaugural feast did its sponsors no good whatever.
Superior new material for orthodontic work is another result of research. He strolled back to the door whistling softly hands still clasped behind him. Her eyes were glazed as if she didn’t hear or even see him.
She found herself able to sing any role and any song which struck her fancy. In most cases these soils are taken up as liquids through capillary action. That doctrine has been accepted by many but has it produced good results.
A smile pulled at the lower strip of adhesive tape. The orchestra was obviously on its mettle and it played most responsively. Then he would realize they were really things that only he himself could think.
Space charge influences will also decrease at increased voltages. She served for a number of years without pay beyond her travel and maintenance. His superiors had also preached this saying it was the way for eternal honor.
In tradition and in poetry the marriage bed is a place of unity and harmony. He injected more vitality into the score than it has revealed in many years. Somehow we old timers never figured we would ever retire.
Sprouted grains and seeds are used in salads and dishes such as chop suey. Its sphere is that of royal courts dynastic quarrels and vaulting ambitions. Not without good reason has the anatomical been called jocular journalese.
Very peculiar retribution indeed seems to overtake such jokers. His portrayal of an edgy head in the clouds artist is virtually flawless. The theory the idea behind our design is modular units or panelization.
You think somebody is going to stand up in the audience and make guilty faces. His name became synonymous with cold blooded cruelty. One of the most common of camp maladies was diarrhoea.
No group of officers came in for more spirited denunciation than the doctors. We seemed to be witnessing the population explosion right in our own backyards. Being based on so few events these results are of dubious validity.
For me it has more of both elements than the majority of its competitors. Adults take a long time to convince and you are thwarted if you try to push. You could also say that in these pamphlets is a relieving quality of maturity.
But considered within technical astronomy a different pattern can be traced. Neither are beds thanks to air mattresses and sleeping bags. Rabies were cured or prevented by madstones which the pioneer wore or carried.
Like a fair number of bootleggers he disliked alcohol. How much and how many profits could a majority take out of the losses of a few. And he was handsome despite the long thin scar that slanted across his cheek.
She encouraged her children to make their own halloween costumes. His first glimpse of the ranch house across the brushy swells told him nothing. They were shown how to advance against an enemy outpost atop a cleared ridge.
Have a test run on the family first to be sure timing and seasoning are right. As coauthors we presented our new book to the haughty audience. To keep ’em scattered somewhat and yet herd ’em was called loose herdin’.
We have also seen the power of faith at work among us. Last year’s gas shortage caused steep price increases. To use these new ways in daily life is the last step.
We now generalize these ideas for general binomial experiments. An area sheltered from strong winds may be highly desirable for recreation use. We would lose our export markets and deny ourselves the imports we need.
A cardboard pattern cut to fit inside holder will help to prevent warping. These were thought to represent regenerating fibers. The wrinkled mouth laughed revealing astonishingly strong white teeth.
It seems that open season upon veterans’ hospitalization is once more upon us. Why don’t they tell me themselves if it bothers them. Cory attacked the project with extra determination.
Now you know she could’ve but she isn’t that kind of girl. It had assumed the terrifying inertia of inanimate matter. It would be well to show the populace how we deal with adulterers.
In an ideological argument the participants tend to thump the table. The narrow fringe of sadness that ran around it only emphasized the pleasure. Worse his present crew included five men who had sailed with him before.
Her face seemed to float in an implausibly bright shaft of sunlight. Sometimes soldiers wrote letters while bullets were whizzing about their heads. The thought came back the one nagging at him these past four days.
Manual leveling requires an appropriate display of the accelerometer outputs. The other patrons were taxi drivers and art students and small shopkeepers. State numbering laws differ from each other in many ways.
But the information on the dynamics of population was often quite misleading. We will achieve a more vivid sense of what it is by realizing what it is not. Originals are not necessarily good and adaptations are not necessarily bad.
The odor here was more powerful than that which surrounded the town aborigines. She drank greedily and murmured thank you as he lowered her head. In my place you’d follow such advice as you give me.
I took her word for it but is she really going with you. Sometimes although by no means always these are indeed alkaline. Mike was of legal age and presumed able to defend himself in the clinches.
As we observe moral law and physical law they appear as being inevitable. He was busy he said in having someone submit to a monkey gland operation. Well i guess i ought to dust out that desk anyhow.
Our hypothetical other bum who killed him would have turned out his pockets. There should be no reason to misinterpret or ignore the intent of this letter. No amount of ballyhoo will cover up the sordid facts.
We may say of some unfortunates that they were never young. The word also made him feel hate sincere hate for those so labeled. Impressions often appear in a symbolic form and cannot be taken at face value.
There are several sources of evidence on the micrometeorite environment. Ice baths electric shocks lashings wild dogs testicle crushers. Caution continuous administration is not recommended for lactating cows.
Ideally he knew it should be preceded by concrete progress at lower levels. The decking is quarter inch mahogany marine plywood. The battery median grade equivalent was used in data analysis in this study.
There are optimal humidity requirements for various agents when airborne. It was the story of the rhinoceros fight all over again. Soil redeposition is evaluated by washing clean swatches with the dirty ones.
What a discussion can ensue when the title of this type of song is in question. Meanwhile fishermen took advantage of them to pull up whoppers. The season between spring and summer belongs to life in its carefree aspect.
Opaque cantaloupe and transparent wood brown were used. Wooded stream valleys in the folds of earth would be saved. He didn’t tell her the truth he now freely admitted to himself.
Now the problem is presented piecemeal and sometimes contradictorily. Questions came to me from all sides about my world citizenship activities. He did not however settle back into acquiescence with things as they were.
For roast insert meat thermometer diagonally so it does not rest on bone. Again the analyticity of the two curves guarantees that such intervals exist. Lips pursed mournfully he stared down at its crazily sagging left side.
During one reading an image appeared of a prisoner in irons. Is there any word you would like to offer in your own defense. Not very well behaved is she to run out on a play mate.
We did not accept the diagnosis at once but gradually we are coming to. Another brand of indefinite reference arises out of the use of the double verb. The failure to keep these two usages distinct presents hazards to the reader.
If you destroy confidence in banks you do something to the economy he said. New self deceiving rags are hurriedly tossed on the too naked bones. Somewhere birds were sweetly calling were answered.
This process is especially difficult since gyro drifting is typically random. Keep your seats boys i just want to put some finishing touches on this thing. The world is constantly changing what was new yesterday is obsolescent today.
Although my shot killed his horse he rolled off the bale on top of me. It latches when you close it so stay as long as you like. Other interpretations present the music as an essentially intimate creation.
She had no way of knowing in advance whether an opportunity for murder existed. No more startling contrast to a system of sullen satellites could be imagined. They make gin saws and deal in parts supplies and some used gin machinery.
A chaw of tobacco put on an open wound was both antiseptic and healing. No antigen was detectable in certain dark spherical areas in most cells. It was like finally getting into one’s own nightmares to punish one’s dreams.
One species of ambiguity tries to baffle by interweaving repetition. Though brief it has a sharp dramatic edge and great poignancy. The continuing modernization of these forces is a costly but necessary process.
Ran away on a black night with a lawful wedded man. If you use parking attendants can they be replaced by automatic parking gates. A monstrous shadow fell across the illuminated wall distorted and indefinable.
It will accommodate firing rates as low as half a gallon an hour. One of these is the solidarity and the confidential relationship of marriage. Other morphological physical and optical property values are also given.
She took it grudgingly her dark eyes baleful as they met his. The highest rated non supervisory engineering title is research engineer. What had been the ambassador’s suite was now jagged walls of blackened brick.
He is not talking in the main about probabilities risks and danger in general. A kerosene shampoo seems a heroic treatment but it did the job. They should live in modest circumstances avoiding all conspicuous consumption.
Measured performance characteristics for this experimental tube will be listed. One of the most desirable features for a park are beautiful views or scenery. The smell is sexual but so powerfully so that a civilized nose must deny it.
Death reminds man of his sin but it reminds him also of his transience. He merely said any good decorator these days can make you a tasteful home. Thinking the evidence insufficient to get a conviction he later released him.
Conservatism and traditionalism seem implied by what has just been said. Thirty five military and civilian students received laboratory training. They did not know who they were or know their own worth.
Pretty girls among them with blonde hair and pert faces. He stared at the far morning expecting a pendulum to swing across the horizon. Also make sure thermometer does not touch the revolving spit or hit the coals.
If it ever got behind me the beep turned to a buzz. We flew in rickety planes so overloaded that we wondered why they didn’t crash. Tetanus could be avoided by pouring warm turpentine over a wound.
So if anybody solicits by phone make sure you mail the dough to the above. But they would reconsider it they assured him if he would rewrite it. Clapping spurs to the bronc he set off at a sharp canter with growing alarm.
She took it with her wherever she went she chose it. The old woman arose stiffly and led me to a clearing where a small hut stood. There was no confirmation of such massive assaults from independent sources.
The walls bulged the floor trembled the windowpanes rattled. Sewing brings numbness writes what makes my hands numb when sewing. For sweet sour sauce cook onion in oil until soft.
Running around in the moonlight almost naked and slugging a man with a rock. He remembered the last time he had eaten actual eggs from an actual pan. The simplest kind of separate system uses a single self contained unit.
You could burn down this whole mountainside with a fire that size. There are many such competently anonymous performances among the earlier poems. If a concessionaire runs the cafeteria keep an eye out for quality and price.
Ralph prepared red snapper with fresh lemon sauce for dinner. This staff deserves a lot of credit working down here under real obstacles. Tiny bodies dropped onto a dry leaf made a pile as big as a small apple.
No chemical fertilizers and poisonous insecticides and fungicides are used. The total of these three volumes is the final combustion chamber volume. We were off the road gleaming barbed wire pulling taut.
Yet it exists and has an objective reality which can be experienced and known. They inhabit a secret world centered on go codes and gold phones. Latest models serve hot meals at reasonable prices and at a profit to you.
Both eventualities are possible logically but practically they are impossible. The system may break down soon so save your files frequently. But why is it necessary to reproduce the retinal image within the brain.
He daydreamed on the rock while she swam and splashed around. Bathing the itching parts with kerosene gave relief and also killed the pests. There is little doubt that the students benefit from vocational education.
The gunman nodded slipping the picture into his breast pocket saying nothing. They were not yet prepared to accept it as irremediable. With any luck at all he could easily find a flowerpot.
Evidence that other sources of financing are unavailable must be provided. We send shovels cement nails and corrugated iron for roofs. The avocado should have a give to it as you hold it when it is ripe.
Prompted by a guilty urge he had disobeyed the order of a man he respected. He will say that our country is even now a homogeneous community. This girl soon drops the bourgeoisie psychiatrist who disapproves of her life.
He slipped outside hugging the walls of buildings and dodging into doorways. And men also used vacuum cleaners in both rooms sucking dust up once more. The dimensions of these waves dwarf all our usual standards of measurement.
Promptly at seven he would clatter out of the court with twelve in the tallyho. Now a distinguished old man called on nine divinities to come and join us. He hoped they would put in somewhere way way down in the earth.
The population can thereby replenish itself and actually grow larger. They offered no opinions volunteered nothing betrayed no emotions. These needs usually concern the reduction of guilt and some relief of tension.
Microorganisms are often responsible for the rapid spoilage of foods. One of the problems associated with the expressway stems from the basic idea. Only incomplete imperfect things move towards what they lack.
We’ll both be blowing town tomorrow so we won’t be moving in on you. It required an energy he no longer possessed to be satirical about his father. No question ruffles him or causes him to hesitate.
Flaxseed poultices and mustard plasters still are used by some persons. All that time rifle barrels were pointing unwaveringly at his head and body. Steam baths writes do steam baths have any health value.
In time she presents her aristocratic husband with a coal black child. The straight line would symbolize its uniqueness the circle of universality. Even then if she took one step forward he could catch her.
Would have been easy to identify as opium by its odor. This is a problem that goes considerably beyond questions of salary and tenure. Quince seed gum is the main ingredient in wave setting lotions.
Half slyly he enjoyed seeing her stoop to lift the things. We hope that he will execute it in a manner that will entitle him to credit. A site may also be attractive just through the beauty of its trees and shrubs.
Her face turned pink with pleasure and a smothered cough. Would a camera club be useful in taking pictures pertinent to plant safety. Family loyalties and cooperative work have been unbroken for generations.
Would a blue feather in a man’s hat make him happy all day. He reached once more into the carpet bag and brought up a package of wieners. There was a grunt curiously inarticulate like that of an animal in pain.
Insulate weatherstrip double glaze to the maximum. He wanted to show the town what happened to anyone who tried to start trouble. Ants carry the seeds so better be sure that there are no ant hills nearby.
You should firmly insist that no bobby pins or hair pins be worn in the water. In some measure they depend upon the structure of individual personality. We congratulate the entire membership on its record of good legislation.
On unoccupied roadway the bottle shattered into a small amber flash. The name fell with lazy affectionate remembrance from her lips. When he finally did he had to duck his head quickly away as the pitch came in.
Afraid you’ll lose your job if you don’t keep your mouth shut. But to the infuriation of scientists for no known reason not all of them did. Meanwhile the enemy will capitalize on our fears if he can.
Further it has its work cut out stopping anarchy where it is now garrisoned. It has multiple implications and possible headaches for your marketing program. One more muddleheaded play like that one and they’d be leading him away.
Traffic frequently has failed to measure up to engineers’ rosy estimates. Ambiguity arises when the pronoun it carries a twofold reference. However the aircraft which we have today are tied to large soft airfields.
We know that actors can learn to portray a wide variety of character roles. The ward was a small one four beds kept reserved for female alcoholics. The sculptor looked at him bugeyed and amazed angry.
But to continue to divorce advanced students from reality is inexcusable. But this doesn’t detract from its merit as an interesting if not great film. Then may i ask where these muddy foot prints came from.
In the lighted interior he saw other men and women struggling into their wraps. Her debut over perhaps the earlier scenes will emerge equally fine. He straightened up alert now as the buffalo hunter came closer.
The crisis later on when debts seemed about to overwhelm me. Something else distracted him yet there was no sound only tomblike silence. Let the orthodontist decide the proper time to start treatment he urges.
Forty seven states assign or provide vehicles for employees on state business. This he added brought about petty jealousies and petty personal grievances. It is possible to make a few generalizations about the six giants themselves.
This has been attributed to helium film flow in the vapor pressure thermometer. As his feet slowed he felt ashamed of the panic and resolved to make a stand. Perhaps one bored holes in the stone with some kind of an electric gadget.
The keelson made of two three inch widths is next installed. Cleaned cloth must be protected against the redeposition of dispersed soil. Brief snips of actual events were shown parades dances street scenes.
His black hat with its wide brim high crown and fur trim rode high. Get copper or earthenware mugs that keep beer chilled or soup hot. Two clotted balls the color of mucus rolled between fiery lids.
Pretend ham make criss cross gashes on one side of skinless frankfurters. The platform accelerometers must be slightly modified for this procedure. They are not true because scientists or prophets say they are true.
The batting average of one success out of seven increased to one out of three. And his relatively small hands and feet gave him an almost delicate appearance. But from the start they had two important ingredients sincerity and realism.
From it spokes of order and degree led to the outward rim of the common man. Resolved that the anti slavery sentiment is becoming ripe for resolute action. But if she wasn’t interested she’d just go back to the same life she’d left.
He looked over at him lying there asleep and he felt a wave of revulsion. In the winter hibachi in the kitchen or grill over the logs of the fireplace. Higher toll rates also are helping boost revenues.
He would offer no theory to account for her murder. He drove sensual patterns off carefully shaving his long upper lip. More often these offices are restricted to the gathering of empirical data.
In these damp circumstances he was an odds on bet to develop pneumonia. The problem of solidarity and morale again involves the concept of values. Maybe you and me will girlie but these two ain’t goin’ nowhere.
Was it a hysterical release from the long strain of vigilance of those weeks. Both the conditions and the complicity are documented in considerable detail. Of particular importance is the study of the actions of drugs in this respect.
Advantages a farm provides a wholesome and healthful environment for children. His eyes were dark fluid fearful and he gave a sigh as my knife went in. It does not indicate loose management ineffective controls or poor policy.
These curves were derived by an analysis of extensive skywave measurement data. Don’t they still call you junior as though you’re about ten years old. Fall slowly forward onto the hands and let the body down to rest on the floor.
Except for those minutes in her room he had lost touch with her as a reality. Replace it with the statue of one or another of the world’s famous dictators. The need for reupholstering redecorating repainting becomes more infrequent.
The beatniks crave a sexual experience in which their whole being participates. So somebody else knew what would happen to her father’s money if she died. If the other pilots were worried they did not show it.
Our first necessity at the very outset of war is post attack reconnaissance. To some extent predispositions are shaped by exposure to group environments. In earlier years the preservation of food was essentially related to survival.
Was she just naturally sloppy about everything but her physical appearance. Sometimes strong stress serves to focus an important secondary relationship. Hired hard lackeys of the warmongering capitalists.
He holds that goodness and badness lie in feelings of approval or disapproval. He murmured to himself with firmness no surrender. We scour literature for them here we find stored the wisdom of great minds.
And these hardy travelers are not unappreciated today. The revolution now under way in materials handling makes this much easier. With those paintings of big constructions crashing down he felt he could stop.
He didn’t figure her at all and if he found out a woman it’d be bad. It offered to surrender its right to exclusive trade but asked an indemnity. He may not rise to the heights but he can get by and eventually be retired.
Her position covers a number of daily tasks common to any social director. A flash illumined the trees as a crooked bolt twigged in several directions. Biological warfare is considered to be primarily a strategic weapon.
They moved toward the skiffs with shocking eagerness elbowing and shoving. Radio reception is cut down by the shielding necessary to keep out radiation. Then he fled not waiting to see if she minded him or took notice of his cry.
Avowed atheists or freethinkers are so rare as to be a curiosity. This big flexible voice with uncommon range has been superbly disciplined. She greeted her husband’s colleagues with smiling politeness offering nothing.
Come sit he repeated motioning to the piled hay bags over the pig leavings. Before that we lumber dealers were working almost single handed on the problem. She smiled and teeth gleamed in her beautifully modeled olive face.
Meanwhile three great terrible forces were coagulating and crystallizing. Residential associations struggle to insulate themselves against intrusions. You’d think her stomach would’ve got used to it in three weeks.
Not good looking but self confident and wise so that it made her attractive. Roleplaying used for analysis follows these general steps leading to training. To prepare mustard cream blend mustard with enough water to make a thin paste.
After another long pause he asked how many people know who they are. Within a system however the autonomy of each member library is preserved. The old shop adage still holds a good mechanic is usually a bad boss.
Beer generally fermented from barley is an old alcoholic beverage. When you’re less fatigued things just naturally look brighter. Quite often honeybees form a majority on the willow catkins.
They were both walking towards each other unhurried. He must rearrange matters so that two performers do not bump into each other. Wet also were the marine’s fatigues and the face had an oily film.
They weren’t as well paid as they should have been. It could take place tomorrow night or it might occur months from now. When we left washington his son tad was ill and mrs lincoln hysterical.
The long and ever increasing column of sportsmen is now moving into a new era. The bacteria formed typical activated sludge floc. The filtered air benefits allergies asthma sinus hay fever.
Are planning and strategy development emphasized sufficiently in your company. His talk turns to what he calls the mess or sometimes this buzzing confusion. Conceivably the submarine defense problem can be solved by sufficient forces.
Two women who had been chattering like parrots were struck dumb. In news items a man is less often shot in the body or head than in the suburbs. This is going to be a language lesson and you can master it in a few minutes.
He doesn’t want her to look frowningly at him or speak to him angrily. But briefly the topping configuration must be examined for its inferences. A vigorous program existed in skiing skating sports and overnight hiking.
This is the second year your mother’s donated my fishing tackle to the bazaar. When did women begin to assert themselves sexually. The good man chortled appreciatively and decided the trip was worth his time.
This explains some group ends and provides a justification of their primacy. Computers are being used to keep branch inventories at more workable levels. Large injection molded letters are also available for sign installations.
The bristles are soft enough to massage the gums and not scratch the enamel. There are canoes ideal for fishing in protected waters or for camping trips. They’ll move around that rock all day following the shade.
It takes a great deal of sophisticated thought to get the impact of this fact. His problem concerns longitudes latitudes and angular velocities. Did you know he is advertising his ham radio equipment for sale this weekend.
The wheel of social life spun around the royal or aristocratic centre. But this esoteric doctrine was lost in the shuffle to acquire special powers. We do not arrive at spatial images by means of the sense of touch by itself.
There are people who travel long distances to assure my continued existence. The same shelter could be built into an embankment or below ground level. Everything in the final analysis reduced itself to sexual symbolism.
Is a relaxed home atmosphere enough to help her outgrow these traits. The elementary school child grows gradually in his ability to work in groups. We’re lost and burning up already she bit out tensely.
Probably around midnight give or take an hour either way. She saw me and sat down beside me three feet away. But i’m so sunburned that every move i make is agony.
Closer still regular barricades of barbed wire hung on timber supports. Cattle which died from them winter storms were referred to as the winter kill. Hiring the wife for one’s company may win her tax aided retirement income.
Paperweight may be personalized on back while clay is leather hard. Blowers should be operated periodically on a regular schedule. There is definitely some ligament damage in his knee.
Another stock vaudeville gag ran mother is home sick in bed with the doctor. He enlisted a staff of loyal experts and of many zealous volunteers. We often say of a person that he looks young for his age or old for his age.
It cost us a hundred thousand dollars and thirty days lost time to fix them. Or borrow some money from someone and go home by bus. Almonds and pistachio nuts are not so high in oil but are rich in protein.
Castor oil made from castorbeans has gone out of style as a medicine. He went on to personal bequests a list of names largely unknown to him. This is no assignment for a frivolous girl she assures him.
The stepmother almost without exception has been presented as a cruel ogress. You’re not living up to your own principles she told my discouraged people. He’d had no idea how unhappy his sweet peach had been.
Women didn’t use white face powder nowadays he recalled. I always say you’ve got a wonderful husband miss margaret. It is one of the rare public ventures here on which nearly everyone is agreed.
The larvae kept warm by the queen are full grown in about ten days. She always did before and showed the utmost confidence in whatever we advised. As he had longed to be he became the echo of a saga.
This viscosity of the material in the drops is of course not negligible. High so it only bounce harmlessly but loudly off a car’s steel roof. Maybe it’s taking longer to get things squared away than the bankers expected.
Can your insurance company aid you in reducing administrative costs. One can only speak of what is in front of him and that now is simply the mess. Push ups push ups are essential but few have the strength for them at first.
In simpler terms it amounts to pointing the platform in the proper direction. From this motor pool personnel develop other meaningful and related data. This one came a bit high at thirty thousand or more.
Most of our aid will go to those nearing self sufficiency. The shock therapies act likewise on the hypothalamic balance. The knifelike pain in his groin nearly brought him down again.
To be passive to be girlishly shy was palpably absurd. Look for these features which may mean you can save duplicate coverage. When he left she knew she would never see him again.
Under this law annual grants are given to systems in substantial amounts. But our stumping tour of the south wasn’t all misery. My beloved ward my perennial gadfly said the whining voice.
His sinuous melody is a sort of naive transcendence of all experience. Anybody carrying anything that might hide a rifle. Woe betide the interviewee if he answered vaguely.
Program note reads as follows take hands this urgent visage beckons us. In either case they do not appreciate the private detective’s zeal. X ray films of the vertebral column showed progression of the demineralization.
Microscopically there was emphysema fibrosis and vascular congestion. In a new house generous roof overhangs are a logical and effective solution. Are you utilizing vending machine proceeds to help pay for your program.
This is what necessitates the nonsystematic character of his astronomy. Individual human strength is needed to pit against an inhuman condition. Knows the score with a supreme effort he broke it off.
Why couldn’t they have dumped him off on someone else. His successors have adopted the opposite alternative. Crooked overlapping twisted or widely spaced teeth.
His artistic accomplishments guaranteed him entry into any social gathering. They played crack the whip a few minutes without mishap.